OpenStudy (anonymous):

The coordinates below represent two linear equations. How many solutions does this system of equations have? Line 1 x y –6 3 3 6 Line 2 x y –3 1 3 3 A. 0 B. exactly 1 C. exactly 2 D. infinitely many

OpenStudy (anonymous):

ok i have re done this twice ow and i am sure it is either a or c

OpenStudy (anonymous):

i am going to recheck AGAIN now

OpenStudy (anonymous):

ok i have gotten a as my final answer

OpenStudy (anonymous):

@livias.random what did u get?

OpenStudy (anonymous):

m = (3-6)/(-6-3) m = -3/-9 m = 1/3 (y-3) = 1/3(x-(-6)) y-3 = 1/3x + 6/3 y = 1/3x + 5 - m = (1-3)/((-3)-3) m = -2/-6 m = 1/3 (y-1) = 1/3(x-(-3)) y-1 = 1/3x + 3/3 y = 1/3x + 2 - 1/3x + 2 = 1/3x+5 2 = 5 This is obviously not true so for no values of x will this work 0 solutions A

OpenStudy (anonymous):

omg yay! i got one right!